# Modeling Annotator Disagreement with Demographic-Aware Experts and Synthetic Perspectives

### v1

@@ 1 Introduction
Substantial disagreement among annotators is common in many subjective annotation tasks, including the classification of online comment toxicity, misinformation labeling, or evaluation of politeness and other communication quality. These disagreements often reflect meaningful differences in perspective rooted in social, cultural, or demographic backgrounds, rather than just random noise Holland and Quinn (1987); Larimore et al. (2021); Sap et al. (2022). For instance, a comment labeled as toxic by younger users might be seen as benign by older ones, reflecting different sensitivities to certain language or topics. Modeling subjective perspectives with annotator disagreements is critical for building systems that reflect and reason over diverse perspectives. However, many existing approaches treat disagreement as noise, collapsing annotations into a single label, which can marginalize minoritized perspectives and lead to biased models Prabhakaran et al. (2021). Several works have treated disagreement as a signal rather than noise Dawid and Skene (1979), addressing it by weighting or filtering items based on disagreement, or learning from the distribution of annotations Uma et al. (2022). Among these, distributional modeling methods have become increasingly prominent, especially in NLP tasks where annotator disagreements reflect socially grounded variations Mostafazadeh Davani et al. (2022); Fleisig et al. (2024); Gordon et al. (2022); Wan et al. (2023). While promising, these methods often lack the inductive biases needed to capture structured group-level variation. Without explicit modeling of shared patterns among similar annotators, models risk underrepresenting marginalized groups.
We introduce DeM-MoE, a Demographic-Aware Mixture of Experts model that learns to represent subjective judgments by routing inputs to expert subnetworks based on annotator demographics. This design introduces an inductive bias: similar annotators may reason about inputs in similar ways. It encourages specialization and improves representation of underrepresented groups. DeM-MoE outperforms strong baselines and state of the art systems across multiple datasets and demographic groups. Beyond architectural innovation, we address modeling disagreement in low-data settings with sparse demographic coverage. We propose using LLMs to generate synthetic annotations via zero-shot prompting and explore training strategies that blend real and synthetic data. Our findings show that the optimal strategies depend on dataset structure. These results highlight how DeM-MoE, combined with strategic data augmentation, can effectively model viewpoint diversity even in low-resource scenarios.
We make the following contributions: 1) We propose DeM-MoE, a modular architecture that uses demographic-aware routing to capture structured variation in annotation behavior across demographic groups. 2) We evaluate the alignment of zero-shot LLM-generated annotations with human ratings on subjective tasks. 3) We present a framework with data generation methods and training strategies to incorporate synthetic data, showing that the configuration’s effectiveness depends on dataset structure. All data and code are available  [^1] http://anonymized.
@@ 2 Related Work
Modeling Individual & Group Preferences in Recommendation Systems. Modeling individual and group preferences is well-studied in recommendation systems. Group methods often aggregate member preferences to make collective decisions, using centrality-aware representations Yin et al. (2022) or hierarchical attention and bi-level frameworks to disentangle individual and group representations Xu et al. (2024); Liang et al. (2022); Wang et al. (2024b). MoE architectures are shown to be effective as well, as they capture multifaceted user interests Liu et al. (2024), model user-item interactions Zhao et al. (2020), and support multi-task personalization Kong et al. (2024). However, no prior work in recommendation or NLP uses MoEs to model both individual and group preferences as we do. While both model variation across individuals and groups, the nature and objectives differ. In recommendation, preferences are idiosyncratic, with weak demographic variation, with the aim of combining diverse signals into a unified group preference. Our task focuses on socio-demographic regularities: annotators from similar groups may evaluate content in similar ways. Instead of ignoring or collapsing cases of disagreement, our approach models and preserves these differences to more faithfully represent diverse perspectives.
@@
Modeling Annotator Disagreement in NLP. Disagreement among annotators is common in subjective NLP tasks like toxicity classification, misinformation detection, and stance analysis. Traditional methods treat disagreements as noise by using majority voting or averaging to create a single ‘gold’ label per instance, which can obscure meaningful variation from underrepresented or minoritized groups Prabhakaran et al. (2021). Alternatives include early work that measured or filtered disagreement to improve data quality Aroyo and Welty (2015); Reidsma and op den Akker (2008); Klebanov and Beigman (2014), and more recent approaches that learn from disagreement directly: a) Uncertainty-based methods, which weight examples by annotation variability Plank et al. (2014); b) Distributional and multi-task models, which use label distributions or treat annotators as tasks Mostafazadeh Davani et al. (2022); c) Annotator modeling frameworks, which predict individual labels using shared encoders and per-annotator heads or embeddings Fleisig et al. (2024); Gordon et al. (2022); Wan et al. (2023). Recent work emphasizes the role of demographics in annotation behavior. Wan et al. (2023) model demographics to predict disagreement, while Gordon et al. (2022) use a jury learning approach to estimate group verdicts. However, most models treat demographic features as input to shared encoders, without architectural diversity or inductive biases to model group-specific reasoning. As a result, they capture individual variation but struggle with group-level differences.
Dataset #Inst #Ann Avg/Inst IAA ($\alpha$) Entropy
Safety 350 123 123.0 0.241 0.742
Offensiveness 1,500 262 8.69 0.287 1.212
Patient Centered Comm. 2,230 589 3.33 0.287 1.492
Politeness 3,718 506 6.74 0.440 1.395
Toxicity 107,620 17,172 4.74 0.272 1.070
@@
Table 1: Dataset statistics. “#Inst” = number of instances, “#Ann” = annotators, “Avg/Inst” = avg. annotators per instance, “IAA” = Krippendorff’s $\alpha$, “Entropy” = mean entropy of annotator ratings per instance.
@@ 3 Data
We evaluate our approach on five datasets with diverse tasks, demographic compositions, and levels of annotator diversity and disagreement (Table 1); full details are in Appendix A.1.1. The Toxicity dataset of Kumar et al. (2021) has annotators rate the level of toxicity for social media content from multiple sources; it is the largest dataset, with the broadest demographic coverage (2,535 combinations) and relatively low agreement (Krippendorff’s $\alpha$=0.27), suggesting that annotators could disagree systematically. Ridge regression using demographic group features shows strong demographic signal, especially for age (2.06) and education (1.11). The Safety dataset (Aroyo et al., 2023) asks annotators to rate the how safe an AI response is; the data includes dense annotations (123 per instance) across 48 demographic combinations, but shows low inter-annotator agreement ($\alpha$=0.24), likely due to the complexity of perceived harmfulness judgments. However, demographic signal is comparatively weak (gender: 0.11, age: 0.12). The Popquorn dataset Pei and Jurgens (2023) includes subsets for Politeness and Offensiveness. The Politeness dataset shows the highest agreement ($\alpha$=0.44) and lower label entropy, indicating greater consistency in judgments of email politeness. The dataset has substantial demographic signal, particularly for race (1.43) and education (0.84). Offensiveness lies between these extremes in terms of demographic diversity, agreement, and variability. The dataset has strong demographic signal, especially for age (1.35) and occupation (1.08). Patient Centered Communication (PCC) rates the speech of doctors speaking to patients along multiple attributes. Details on how PCC is collected are in Appendix A.1.2. PCC presents a challenging modeling task: it has sparse annotations per instance (3.33), with the highest uncertainty or ambiguity of labels (entropy), due to the nuanced and personal nature of patient–provider interactions. PCC also shows the strongest demographic signal overall, particularly for age (2.38) and gender (1.54). These datasets vary in the balance between idiosyncratic and systematic demographic disagreement.
@@ 4 DeM-MoE for Modeling Disagreement
Recent work in modeling annotator disagreement has moved beyond majority-vote aggregation to approaches that predict annotator-specific labels Mostafazadeh Davani et al. (2022); Fleisig et al. (2024); Gordon et al. (2022). While many incorporate annotator identity or demographic information, they typically do so by concatenating demographic data with text input Fleisig et al. (2024); Gordon et al. (2022); Wan et al. (2023), or by learning annotator-specific perspectives implicitly through embeddings Deng et al. (2023). By modeling all annotator variation through a single network, these methods do not explicitly enable the model to specialize in distinct judgment patterns across demographic groups. As a result, the model’s ability to systematically represent shared behaviors within demographic subgroups is limited.
We propose a novel approach, DeM-MoE (Demographic-aware mixture of experts), based on Mixture of Experts (MoE) (Fig. 5) Shazeer et al. (2017), which naturally supports modular specialization (different experts learn distinct annotation patterns linked to demographic groups) and selective routing (inputs are dynamically directed to relevant experts based on annotator demographics). Our architecture encodes inductive bias: annotators from similar demographic groups may share systematic ways of judging texts. We address a key gap in prior work: the lack of structured inductive bias. While large networks can learn subgroup variation, they may not do so in a structured or interpretable way. Additionally, they may overfit to dominant groups without an architectural signal promoting subgroup differentiation. Our model makes this inductive bias more robust and interpretable, especially under data imbalance or sparse subgroup representation. It has three components: 1) learned annotator and demographic embeddings; 2) expert selection and dynamic routing; and 3) expert load balance and specialization via a weighted training loss. We discuss each next.
@@
Expert Selection. To promote expert specialization aligned with demographic-group preferences, we build a pool of $n$ experts ($n=$ number of demographic groups). A lightweight gate maps the input vector to expert scores: $\mathbf{s}=W_{s}\,\mathbf{x}+\mathbf{b},\qquad\mathbf{p}=\mathrm{softmax}(\mathbf{s})$. We hard-select the top $k$ experts ($k\in\{2,3\}$): $\mathcal{I}_{k}=\operatorname{arg\,topk}(\mathbf{p},k)$. Gating weights $p_{j}$ determine which demographic-aware experts are used. Each expert $f_{j}$ receives the same input; their outputs are mixed sparsely: $\mathbf{h}\;=\;\sum_{j\in\mathcal{I}_{k}}p_{j}\,f_{j}(\mathbf{x})$. A linear regression head predicts the snippet rating. Ratings are z-score normalized to stabilize training. We prefer hard expert selection over soft gating to reinforce specialization.
Expert Load Balance & Specialization. Naive routing often leads to expert load imbalance, routing collapse, or bottlenecks. While some auxiliary losses improve hardware efficiency via uniformity Fedus et al. (2021); Wang et al. (2024a), we focus on expert diversity with a normalized standard deviation loss, allowing roughly even usage, rather than strict uniformity, to support demographic specialization. We also include an orthogonality loss to encourage distinct expert features, and a variance loss to promote diverse routing paths Guo et al. (2025). Our training loss further encodes the inductive bias, where we include regularizations for annotator and demographic embeddings, in addition to the load loss, orthogonality, variance loss, and demographic within-group specialization loss (see Appendix B).
@@ 5.1 Experiment Setup
Models. We evaluate DeM-MoE against five other models. (1) Probabilistic Matrix Factorization (PMF; Salakhutdinov and Mnih, 2007) is a recommender-system approach that assumes the annotator-item interactions are generated from a probability distribution. As PMF does not make use of the text or any annotator information, it serves as a strong baseline for modeling annotator regularity. (2) As a text-only approach, we use ModernBERT Warner et al. (2024) to predict the rating; this approach does not use any annotator or demographic information. We regress on the embeddings to predict annotator ratings. (3) Orlikowski et al. (2025) proposed learning annotator and demographic variation by LoRA fine-tuning LLaMA 3.1-8B using a sociodemographic prompt. Their results improved over text-only models but revealed the model did not benefit from demographics. (4) The Annotation and Annotator Embedding Model (En + Ea model) Deng et al. (2023) is a SOTA system for explicitly modeling both item and annotator-level variance but does not incorporate demographic information. (5) The Jury Learning Model Gordon et al. (2022) is the closest system to ours. Texts are encoded with ModernBERT, concatenated with annotator group demographic embeddings and annotator embeddings, and passed through a combination of cross and deep networks Wang et al. (2021) to predict each annotator’s rating.
Training. Train/dev/test splits are created at the instance level, ensuring no overlap of snippets between the splits while allowing some overlap in annotators between the train and test sets. Offensiveness has 92% overlap. Politeness has 91%. Safety has 85%. Toxicity has 40%. PCC has 87%. This evaluation focuses on how well the model generalizes to new content and unfamiliar annotators. Training details are in Appendix B.1.
@@
Evaluation. Following the evaluation framework from Gordon et al. (2022), we use Mean Absolute Error (MAE) to measure the prediction accuracy at the annotator level.
@@ 5.2 Results
DeM-MoE consistently performs competitively across demographic groups, with particularly strong results on datasets with high annotator disagreement: Toxicity, Offensiveness, and PCC (Figure 1). Our model outperforms all other models in every demographic group on Toxicity, the most demographically diverse data with low annotator agreement. Our model outperforms all other models in every demographic group except for race on Offensiveness, and gender (but it is statistically equivalent to the best model) on PCC. On Politeness, it is statistically tied with the best models for gender and education.
@@
Second, among models that use both annotator identity and demographics, we see a trend that increasing model structure generally benefits performance and supports our hypothesis (H1) that models with stronger inductive biases can better learn regularity in label variation. LLaMA has the least structure, encoding identity and demographics as text and learning preferences using next-token prediction. Similar to Orlikowski et al. (2025), we find that their model struggles to capture demographic variation on datasets with strong demographic signals (Offensiveness, Politeness, PCC). However, it performs best on Safety, where demographic influence is minimal, suggesting that LLaMA excels in text-dominant settings with limited demographic variability, while more structured models like DeM-MoE offer greater robustness across a broader range of labeling conditions. The consistently high performance of DeM-MoE in settings with high demographic signal suggests that expert routing provides a more effective inductive bias than the dense cross-network architecture and a strong capacity to represent fine-grained differences in annotator viewpoints.
Does Experts Align with Demographics? To test whether the model’s inductive biases lead to experts aligning with demographic groups, we analyze two types of specialization: within-group and cross-group. Within-group specialization focuses on diversity among subgroups within a demographic (e.g., different racial identities), reflecting our view that not all perspectives within a group can or should be collapsed into one. We quantify this by computing the mean pairwise KL divergence in expert usage distributions across subgroups for each demographic. To normalize for model capacity, we divide each KL score by $\log(K)$, where $K$ is the number of experts. We find that experts specialize in capturing variation within demographic groups (Figure 2), but the relevant groups vary by dataset—highlighting interactions between the construct being modeled and demographic-specific variation. For example, the model shows the most specialization on Politeness for race, the strongest predictor of rating variance Pei and Jurgens (2023). Heatmaps of expert usage show different experts specializing in subgroup perspectives (Figures 711). DeM-MoE routing provides more granular modeling of demographics meaningfully tied to prediction.
@@
What kind of data influences performance? Our results show that the effectiveness of DeM-MoE is closely tied to the dataset’s properties. Three key factors mediate performance. (1) Low agreement: DeM-MoE performs best in tasks with low IAA (Offensiveness, Toxicity, PCC), where subjective interpretations vary. In such settings, expert specialization provides a clear advantage over models that collapse annotations into a single label. In contrast, tasks with high IAA (Politeness) offer less room for modeling perspective diversity, limiting the relative benefits of DeM-MoE. (2) Demographic predictiveness of ratings: DeM-MoE performs best when demographic attributes are predictive of annotator ratings (e.g., age and gender in PCC). Expert specialization aligns well with this variance, as performance is weaker when demographics are less predictive (Safety). (3) Annotation density: DeM-MoE benefits from having sufficient annotations per demographic profile to support expert learning. Demographic error was correlated (rr=0.40) with the number of annotations for that group; where datasets like Toxicity with many annotations per demographic combination (213) can be modeled more effectively than those like Politeness with lower density (44).

### v3

@@ 1 Introduction
Substantial disagreement among annotators is common in subjective tasks such as toxicity detection, misinformation labeling, and politeness evaluation. These disagreements often reflect meaningful differences in perspective rooted in social, cultural, or demographic backgrounds rather than random noise Holland and Quinn (1987); Larimore et al. (2021); Sap et al. (2022). For example, a comment judged toxic by younger users might seem benign to older ones, reflecting differing sensitivities to language and topics. Modeling such perspective variation is essential for building systems that represent and reason over diverse viewpoints. Yet many approaches treat disagreement as noise, collapsing annotations into a single label and marginalizing minoritized perspectives Prabhakaran et al. (2021). Recent work instead treats disagreement as signal Dawid and Skene (1979), using weighting, filtering, or learning from annotation distributions Uma et al. (2022). Distributional modeling has become prominent in NLP tasks where disagreements reflect socially grounded variation Mostafazadeh Davani et al. (2022); Fleisig et al. (2024); Gordon et al. (2022); Wan et al. (2023). However, such methods often lack inductive biases to capture structured group-level variation, risking underrepresentation of marginalized perspectives.
We introduce DeM-MoE, a Demographic-Aware Mixture of Experts model that learns to represent subjective judgments by routing inputs to expert subnetworks based on annotator demographics. This design introduces inductive bias: similar annotators may reason about inputs similarly, encouraging specialization and improving representation of underrepresented groups. DeM-MoE outperforms strong baselines and SOTA systems across multiple datasets and demographic groups. Beyond architectural innovation, we address modeling disagreement in low-data settings with sparse demographic coverage using LLM-generated synthetic annotations and blended training strategies. Optimal strategies depend on dataset structure, showing how DeM-MoE, combined with strategic data augmentation, effectively models viewpoint diversity even in low-resource scenarios.
We make three contributions: (1) We propose DeM-MoE, a modular architecture with demographic-aware routing to capture structured variation in annotation behavior. (2) We evaluate the alignment of zero-shot LLM-generated annotations with human ratings. (3) We present a framework for incorporating synthetic data, showing that its effectiveness depends on dataset structure. [^1] Data and code are available at http://anonymized.
@@ 2 Related Work
Modeling Individual & Group Preferences in Recommendation Systems. Recommendation systems extensively model individual and group preferences. Group-based methods aggregate member choices into collective decisions, often using centrality-aware representations Yin et al. (2022) or hierarchical attention to separate individual and group signals Xu et al. (2024); Liang et al. (2022); Wang et al. (2024b). Mixture-of-Experts (MoE) architectures further capture multifaceted user interests Liu et al. (2024), model user–item interactions Zhao et al. (2020), and enable multi-task personalization Kong et al. (2024), including user and group modeling Gong et al. (2023); Liu et al. (2025). In contrast, no prior NLP work uses MoEs to jointly model individual and group variation. While our task and recommenders model variation across individuals and groups, the nature and objectives differ. Whereas recommenders target idiosyncratic preferences with weak demographic patterns and aim to merge signals into a unified ranking across items, our task focuses on predicting singular judgments using systematic socio-demographic regularities. Annotators from similar groups may evaluate content in shared ways, and rather than collapsing disagreement, our approach explicitly preserves them to represent diverse perspectives.
@@
Modeling Annotator Disagreement in NLP. Disagreement among annotators is common in subjective NLP tasks like toxicity classification, misinformation detection, and stance analysis. Traditional methods treat disagreements as noise by using majority voting or averaging to create a single "gold" label per instance, which can obscure meaningful variation from underrepresented or minoritized groups Prabhakaran et al. (2021). Alternatives include early work that measured or filtered disagreement to improve data quality Aroyo and Welty (2015); Reidsma and op den Akker (2008); Klebanov and Beigman (2014), and more recent approaches that learn from disagreement directly: a) Uncertainty-based methods, which weight examples by annotation variability Plank et al. (2014); b) Distributional and multi-task models, which use label distributions or treat annotators as tasks Mostafazadeh Davani et al. (2022); c) Annotator modeling frameworks, which predict individual labels using shared encoders and per-annotator heads or embeddings Fleisig et al. (2024); Gordon et al. (2022); Wan et al. (2023). Recent work emphasizes the role of demographics in annotation behavior. Wan et al. (2023) model demographics to predict disagreement, while Gordon et al. (2022) use a jury learning approach to estimate group verdicts. However, most models treat demographic features as input to shared encoders, without architectural diversity or inductive biases to model group-specific reasoning. As a result, they capture individual variation but struggle with group-level differences.
Dataset #Inst #Ann Avg/Inst IAA ($\alpha$) Entropy Max Coef
Safety 350 123 123.0 0.241 0.742 Race (0.559)
Offensiveness 1,500 262 8.69 0.287 1.212 Age (1.351)
Patient Centered Comm. 2,230 589 3.33 0.287 1.492 Age_group (2.528)
Politeness 3,718 506 6.74 0.440 1.395 Race (1.427)
Toxicity 107,620 17,172 4.74 0.272 1.070 Age_range (2.056)
@@
Table 1: Dataset statistics. ‘#Inst” = number of instances, #Ann” = annotators, Avg/Inst” = avg. annotators per instance, IAA” = Krippendorff’s $\alpha$, Entropy” = mean entropy of annotator ratings per instance. “Max Coef” = demographic feature with the largest absolute ridge regression coefficient per dataset.
@@ 3 Data
We evaluate on five datasets spanning diverse tasks, demographic coverage, and levels of disagreement (Table 1; see full details in Appendix A.1.1). The Toxicity dataset (Kumar et al., 2021) is the largest, with the broadest demographic coverage (2,535 combinations) and low agreement ($\alpha$=0.27), suggesting systematic disagreement. Ridge regression predicting annotator ratings using demographics shows strong demographic signal, especially for age (2.06). The Safety dataset (Aroyo et al., 2023) contains dense annotations (123 per instance) across 48 combinations, but has low agreement ($\alpha$=0.24). Demographic signal is weak likely due to the complexity of harmfulness judgments. The Popquorn dataset Pei and Jurgens (2023) includes Politeness and Offensiveness. Politeness shows the highest agreement ($\alpha$=0.44) and low entropy, with strong demographic effects (race). Offensiveness shows notable signal for age and occupation. The Patient Centered Communication (PCC) dataset rates doctor–patient interactions across multiple attributes [^2] Details on how PCC is collected are in Appendix A.1.3. It has sparse annotations (3.3 per instance) and high uncertainty (entropy). PCC also shows the strongest demographic signal, especially for age. These datasets highlight varying balances of idiosyncratic vs. systematic disagreement, underscoring the need for models that preserve demographic variation.
@@ 4 DeM-MoE for Modeling Disagreement
Recent work on annotator disagreement has moved beyond majority-vote aggregation toward models that predict annotator-specific labels Mostafazadeh Davani et al. (2022); Fleisig et al. (2024); Gordon et al. (2022). Many incorporate annotator identity or demographics, but typically by concatenating demographic features with text Fleisig et al. (2024); Gordon et al. (2022); Wan et al. (2023) or by implicitly encoding annotator perspectives through embeddings Deng et al. (2023). Because these systems model all annotators through a single network, they cannot explicitly specialize in distinct judgment patterns across demographic groups. Consequently, their ability to represent shared group behaviors remains limited. Orlikowski et al. (2023) partially addresses this by injecting demographic information into a shared encoder with group-specific layers, but intersectional identities are still modeled through separate, independent components, constraining the model’s capacity to learn intersectional patterns.
We propose a new approach, DeM-MoE (Demographic-aware mixture of experts), based on Mixture of Experts (MoE) (Fig. 5) Shazeer et al. (2017), which naturally supports modular specialization (different experts learn distinct annotation patterns linked to demographic groups) and selective routing (inputs are dynamically directed to relevant experts based on annotator demographics). The input consists of the text snippet (encoded with Modern-BERT), the annotator embedding, and the demographic embeddings, and then all three are concatenated into the MoE input, which is then routed to experts to predict the annotator’s rating. Our architecture encodes inductive bias: annotators from similar demographic groups may share systematic ways of judging texts. We address a key gap in prior work: the lack of structured inductive bias. While large networks can learn subgroup variation, they may not do so in a structured or interpretable way. They may also overfit to dominant groups without an architectural signal promoting subgroup differentiation. Our model makes this inductive bias more robust and interpretable, especially under data imbalance or sparse subgroup representation. Our model’s routing also allows experts to naturally specialize in intersectional subgroups, and learn cross-group and within-group specialization. We discuss the model components next: 1) learned annotator and demographic embeddings; 2) expert selection and dynamic routing; and 3) expert load balance and specialization via a weighted training loss.
@@
Expert Selection. To promote expert specialization aligned with demographic-group preferences, we build a pool of $n$ experts ($n=$ number of demographic groups). A lightweight gate maps the input vector to expert scores: $\mathbf{s}=W_{s}\,\mathbf{x}+\mathbf{b},\mathbf{p}=\mathrm{softmax}(\mathbf{s})$. We hard-select the top $k$ experts ($k\in\{2,3\}$): $\mathcal{I}_{k}=\operatorname{arg\,topk}(\mathbf{p},k)$. Gating weights $p_{j}$ determine which demographic-aware experts are used. Each expert $f_{j}$ receives the same input; their outputs are mixed sparsely: $\mathbf{h}\;=\;\sum_{j\in\mathcal{I}_{k}}p_{j}\,f_{j}(\mathbf{x})$. A linear regression head predicts the snippet rating. Ratings are z-score normalized to stabilize training. We prefer hard expert selection over soft gating to reinforce specialization.
Expert Load Balance & Specialization. Naive routing often leads to expert load imbalance, routing collapse, or bottlenecks. While some auxiliary losses improve hardware efficiency via uniformity Fedus et al. (2021); Wang et al. (2024a), we focus on expert diversity with a normalized standard deviation loss, allowing roughly even usage, rather than strict uniformity, to support demographic specialization. We also include an orthogonality loss to encourage distinct expert features, and a variance loss to promote diverse routing paths Guo et al. (2025). Our training loss further encodes the inductive bias, where we include regularizations for annotator and demographic embeddings, load loss, orthogonality, variance loss, and demographic within-group specialization loss (see Appendix B).
@@ 5.1 Experiment Setup
All models are evaluated to predict individual annotator ratings using the text of each item and, where applicable, annotator or demographic information. Each model differs in its inputs (text only vs. text + annotator/demographics), while the output is a scalar rating per annotator–item pair, except for ModernBERT (which predicts at the snippet level with text input). (1) Probabilistic Matrix Factorization (PMF) (Salakhutdinov and Mnih, 2007) uses only annotator and item IDs, learning latent embeddings for both and modeling their interaction as a probability distribution over ratings, thus capturing annotator regularity without text or demographics. (2) As a text-based baseline, we fine-tune ModernBERT-large (Warner et al., 2024) with a single linear regression head on the CLS embedding to predict the mean snippet rating. (3) LLaMA-3.1-8B-Instruct (Orlikowski et al., 2025) receives text alone in a zero-shot setting or text plus annotator demographic descriptions when LoRA-fine-tuned with sociodemographic prompts. Their results improved over text-only models but revealed the model did not benefit from demographics. (4) Annotation + Annotator Embedding model (En + Ea) (Deng et al., 2023) is a SOTA system for explicitly modeling both item and annotator-level variance but does not incorporate demographic information. (5) Jury Learning (Gordon et al., 2022) uses ModernBERT text embeddings, annotator embeddings, and group-level demographic embeddings, concatenated and passed through cross and deep networks (Wang et al., 2021) to predict annotator ratings; this architecture most closely parallels ours.
Training. Train/dev/test splits are created at the instance level, ensuring no overlap of snippets between the splits while allowing some overlap in annotators between the train and test sets. Offensiveness has 92% overlap, Politeness 91%, Safety 85%, Toxicity 40%, and PCC 87%. This evaluation focuses on how well the model generalizes to new content and unfamiliar annotators. Training details are in Appendix B.1.
@@
Evaluation. We use Mean Absolute Error (MAE) to measure the average absolute different between the model’s predicted rating and actual annotator rating, at the individual levelGordon et al. (2022); Fleisig et al. (2024). For analysis, we aggregate across annotators’ demographic groups.
@@ 5.2 Results
DeM-MoE consistently performs competitively across demographic groups, with particularly strong results on datasets with high annotator disagreement: Toxicity, Offensiveness, and PCC (Figure 1). [^3] We also test best zero-shot performance with LLaMA 3.1-8B-Instruct, but omit it from the figure due to its low performance. Results are reported in Appendix C.2 Our model outperforms all other models in every demographic group on Toxicity, the most demographically diverse data with low annotator agreement. Our model outperforms all other models in every group except for race on Offensiveness, and gender (but statistically equivalent to the best model) on PCC. On Politeness, it is statistically tied with the best models for gender and education.
@@
Second, among models that use both annotator identity and demographics, we see a trend that increasing model structure generally benefits performance and supports our hypothesis (H1) that models with stronger inductive biases can better learn regularity in label variation. LLaMA has the least structure, encoding identity and demographics as text and learning preferences using next-token prediction. Similar to Orlikowski et al. (2025), we find that their model struggles to capture demographic variation on datasets with strong demographic signals (Offensiveness, Politeness, PCC). However, LLaMA performs best on the Safety task, where demographic influence is minimal. This suggests that it excels in text-dominant settings with limited demographic variability (and benefits from having roughly 1M more parameters than our model). In contrast, DeM-MoE is worse and likely underfits in such settings, as the routing function provides limited benefit when individual and group preferences are weak or absent. However, more structured models like DeM-MoE offer greater robustness across a broader range of labeling conditions. The consistently high performance of DeM-MoE in settings with high demographic signal suggests that expert routing provides a more effective inductive bias than the dense cross-network architecture and a strong capacity to represent fine-grained differences in annotator viewpoints.
While larger model capacity might explain performance boost, we observe this is not the case (Tab. 25). Models with far larger parameter counts i.e. LoRA-LLaMA (3.4M) often underperform MoE (2.5M), suggesting that our advantage stems from inductive bias rather than increased capacity.
Do Experts Align with Demographics? To test whether the model’s inductive biases lead to experts aligning with demographic groups, we analyze two types of specialization: within-group and cross-group. Within-group specialization focuses on diversity among subgroups within a demographic (e.g., different racial identities), reflecting our view that not all perspectives within a group can or should be collapsed into one. We quantify this by computing the mean pairwise KL divergence in expert usage distributions across subgroups for each demographic. To normalize for model capacity, we divide each KL score by $\log(K)$, $K$ = the number of experts. We find that experts specialize in capturing variation within demographic groups (Figure 2), but the relevant groups vary by dataset –highlighting interactions between the construct being modeled and demographic-specific variation. The model shows the most specialization on Politeness for race, the strongest predictor of rating variance Pei and Jurgens (2023). Heatmaps of expert usage, aggregated across all instances within each demographic category, show different experts specializing in subgroup perspectives (Figures 711). [^4] The diffuse activation patterns arise because averaging across instances makes multiple experts appear active, reflecting subgroup variability rather than soft gating. MoE routing provides more granular modeling of demographics meaningfully tied to prediction.
@@
What kind of data influences performance? The effectiveness of DeM-MoE is closely tied to the dataset’s properties. Three key factors mediate performance: (1) Low agreement: MoE performs best in tasks with low IAA (Offensiveness, Toxicity, PCC), where subjective interpretations vary. In such settings, expert specialization provides a clear advantage over models that collapse annotations into a single label. In contrast, tasks with high IAA (Politeness) offer less room for modeling perspective diversity, limiting the relative benefits of MoE. (2) Demographic predictiveness of ratings: MoE performs best when demographic attributes are predictive of annotator ratings (e.g., age and gender in PCC). Expert specialization aligns well with this variance, as performance is weaker when demographics are less predictive (Safety). (3) Annotation density: MoE benefits from having sufficient annotations per demographic profile to support expert learning. Demographic error was correlated (rr=0.40) with the number of annotations for that group; where datasets like Toxicity with many annotations per demographic combination (213) can be modeled more effectively than those like Politeness with lower density (44).